OpenStudy (anonymous):

I am doing a related rates problem and I don't know how to do implicit differentiation. Find ds/dt s=4000+60x+0.02x^2

jimthompson5910 (jim_thompson5910):

s=4000+60x+0.02x^2 d/dt[ s ]=d/dt[ 4000+60x+0.02x^2 ] ds/dt =d/dt[ 4000+60x+0.02x^2 ] ds/dt =d/dt[ 4000 ]+d/dt[ 60x ]+d/dt[ 0.02x^2 ] ds/dt =0 +60*d/dt[ x ]+0.02*d/dt[ x^2 ] ds/dt =0 +60*dx/dt+0.02*2x*dx/dt ds/dt = 60*dx/dt+0.04x*dx/dt